OpenStudy (anonymous):

since 3^0=1 what is the value of log 3^1

Parth (parthkohli):

\[\log_{\large3} 1\]For the answer, ask yourself: "What power of 3 results in 1?"

Parth (parthkohli):

Hint: You are already given the power of 3 which results in 1.

OpenStudy (klimenkov):

You just need a definition of the logarithm to answer your question.

OpenStudy (anonymous):

i got that one already. Can you help me with (1/2)^256 please

OpenStudy (anonymous):

i cannot figure out how to enter this in my calculator, but i also would like to know how to do it without my calculator. but for now what would the answer be.

OpenStudy (klimenkov):

\((\frac12)^{256}\) ??

OpenStudy (anonymous):

yes

OpenStudy (klimenkov):

Can you tell me why do you want to find it? It is a very small number. http://www.wolframalpha.com/input/?i=%281%2F2%29^256

OpenStudy (anonymous):

it is a problem that i am working on in my math lab

OpenStudy (anonymous):

exponential form

OpenStudy (klimenkov):

It depends on abilities of your calculator. Click the link I have given you and see what number it is.

OpenStudy (anonymous):

cheesy calculator

OpenStudy (anonymous):

cheap

OpenStudy (klimenkov):

If you dont have an exponent on your calculator, you have to press 0.5x0.5x0.5x...x0.5 256 times. I think it is not worth of it.

OpenStudy (anonymous):

so 0.5x256?

OpenStudy (anonymous):

or 0.5x0.5x0.5x0.5x256?

OpenStudy (anonymous):

128?

OpenStudy (klimenkov):

No. Not 0.5+0.5+...+0.5. You have to multiply \(\frac12\cdot\frac12\cdot\frac12\cdot\frac12\cdot\frac12\cdots\frac12\) 256 times!!

OpenStudy (anonymous):

that would take forever

OpenStudy (anonymous):

i dont have all day. lol

OpenStudy (klimenkov):

Not forever but a long time.

OpenStudy (klimenkov):

Why do you want to find it? May be there is another problem?

OpenStudy (anonymous):

oh gosh, can you help me? do you have a better calculator than me? how do i enter this in a calculator?

OpenStudy (klimenkov):

Did you go through the link I had given you?

OpenStudy (anonymous):

i see no link

OpenStudy (klimenkov):

http://www.wolframalpha.com/input/?i=%281%2F2%29^256

OpenStudy (anonymous):

oh wolfram. i have that, but how would i enter this in wolfram

OpenStudy (klimenkov):

(1/2)^256 it will understand

OpenStudy (anonymous):

k i got 10^-78 is that right

OpenStudy (anonymous):

or is it 8.3636

OpenStudy (anonymous):

i mean 8.636

OpenStudy (klimenkov):

it is \(8.636\cdot10^{-78}\)

OpenStudy (anonymous):

now do i simplify that

OpenStudy (anonymous):

or is that my final answer

OpenStudy (klimenkov):

It is the final answer.

OpenStudy (anonymous):

k hold on let me type that in my math lab and see if it is correct

OpenStudy (anonymous):

nope the answer is y,256

OpenStudy (klimenkov):

What is the first problem you need to solve?

OpenStudy (anonymous):

hold on and i will write everything out

OpenStudy (anonymous):

find the exact value of the logarithm without using a calculator, in exponential notation use fact that if a^u=a^v then u=v. log 1/2^256 y=log 1/2^256 change to exponential form. Use the fact that 1/a=a^-1 to rewrite the left side of the equation. (1/2)^y=256 ____=256.

OpenStudy (anonymous):

what would i put in this blank

OpenStudy (anonymous):

are you still there

OpenStudy (anonymous):

hello

OpenStudy (anonymous):

are you there or not

OpenStudy (klimenkov):

Hm.. I can't understand the question. I think (1/2)^y=256 is an exponential form.